OpenStudy (anonymous):

How many moles of methane are produced when 25.1 moles of carbon dioxide gas react with excess hydrogen gas?

OpenStudy (kellyartigues):

The balanced equation is CO2 + 4 H2 = CH4 + 2 H2O the ratio between CO2 and CH4 is 1 : 1 moles CH4 = 25.1 the ratio between H2 and H2O is 4 : 2 moles H2 = 19.1 x 4 / 2 = 38.2